two vectors each have magnitudes of 18 and 5. what is the minimum and maximum value the resultant of these two vectors can have?
1
Expert's answer
2012-10-09T09:12:38-0400
The maximum resultant can be obtained if the given vectors are directed equally. Then the magnitude of this resultant is

Max = 18 + 5 = 23.

The minimum resultant can be obtained if the given vectors have opposite directions. Then the magnitude of this resultant is

Min = 18 - 5 = 13.